An airline claims that the no-show rate for passengers is less than 5%. In a sample of 420 randomly selected reservations, 19 were no-shows. At   , test the airline’s claim.

State the sample percentage.
Answer #1

P = * = 120 = 0.0452 The sample percentage = p *100 = 0.0452*100 =4.52 % Null and alternative Hypothesis Họ: p= 0.05 vs HA: p